OpenStudy (anonymous):

what is the simplified form of x+2/x^2-3x-10 times x-3/x^2+x-12

OpenStudy (anonymous):

factorize x^2-3x-10 =(x-5)(x+2) and x^2+x-12 =(x-3)(x+4) ( x+2/x^2-3x-10 )( x-3/x^2+x-12) =(x+2)/((x-5)(x+2)) (x-3)/(x-3)(x+4) =1/(x-5) 1/(x+4) =1/((x-5)(x+4))

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

you are welcome :D